1. A 1.22 mol sample of krypton gas at a temperature of 20.0 °C is found to occupy a volume of 23.4 liters. The pressure of this gas sample is  mm Hg.

2. A sample of hydrogen gas at a pressure of 765 mm Hg and a temperature of 24 °C, occupies a volume of 5.39liters. If the gas is cooled at constant pressure to a temperature of 2 °C, the volume of the gas sample will be

3. A 0.532 gram sample of helium gas has a volume of 989 milliliters at a pressure of 2.78 atm. The temperature of the He gas sample is
Answer #1

No of moles of gas (n) = 1.22

Temperature (T) = 20oC = 293K [T(in K) = T (in oC) + 273]

Volume (V) = 23.4 Litres

Universal gas constant (R) = 0.0821 L atm K-1 mol-1

Pressure ( P) = ?

Using Ideal Gas equation,

PV = nRT

P × 23.4 L = 1.22 moles × 0.0821L atm K-1 mol-1 x 293K

P = 1.25 atm

Now 1 atm = 760 mm Hg

1.25 atm = 760 × 1.25 = 950 mm Hg

If you want answer to other questions as well, please post them seperately, preferably 1 at a time